how many grams of sodium phosphate(Na3PO4) are required to make 0.35 liters of a 0.5 M solution?
1
Expert's answer
2020-05-03T15:10:01-0400

0.35*0.5 = 0.175 mol of Na3PO4 - in the solution

M (Na3PO4) = 163.94 g/mol

m (Na3PO4) = 0.175 * 163.94 = 28.6895 g